Deon has 12ml of a 5% solution that has a concentration of 30%. How many ml of 30% solution does he need to add to the 5% solution to make a 20% solution?
Answer by richwmiller(17219)

We have 12 mL of 5% solution
We want to add some 30% solution to make 20% solution
0.3*x+0.05*12=0.2(12+x)
0.3*x+0.6=2.4+0.2x
0.3x-0.2x=2.4-0.6
0.1x=1.8
x=18 mL at 30% (5.4) is added to 12 mL at 5%(0.6) to make 20% of 30 mL(6)
check
0.3*18+0.05*12=0.2(12+18)
5.4+0.6=0.2(30)
6=6
ok
also see
http://www.algebra.com/algebra/homework/word/misc/Miscellaneous_Word_Problems.faq.question.568749.html
with one right answer and one wrong.